A tractor traveling at80m/s on level ground suddenly sees a cow eating grass 200m away in the middle of the farm. If the minimum braking force needed to stop the tractor is 800N. Calculate the mass of the tractor


1
Expert's answer
2021-02-05T11:47:38-0500

Applying work-kinetic energy theorem, we get:


"\Delta KE=W=-F_{br}d,""0-\dfrac{1}{2}mv_i^2=-F_{br}d,""m=\dfrac{2F_{br}d}{v_i^2},""m=\dfrac{2\cdot800\ N\cdot200\ m}{(80\ \dfrac{m}{s})^2}=50\ kg."